OpenStudy (anonymous):

In class of 77 students, there are 5 more girls than twice the number of boys. How many girls are in the class?

OpenStudy (anonymous):

well, we have to make an equation out of it b + g = 77 g = 5 + 2b b +(5 + 2b) = 77 b + 5 + 2b = 77 3b + 5 = 77 3b = 77 - 5 3b = 72 b = 72/3 b = 24 24 + g = 77 g = 77 - 24 g = 53 their are 53 girls

OpenStudy (anonymous):